OpenStudy (anonymous):

\[\[\lim_{x \rightarrow 6}(\frac{ \sqrt[5]{x^2-4}-2 }{ x-6 })\]\]

OpenStudy (anonymous):

Evaluate the following limits

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

HELPP Q~Q

OpenStudy (shamil98):

that looks scary

OpenStudy (anonymous):

@nincompoop

OpenStudy (nincompoop):

how is that a logarithmic problem?

OpenStudy (anonymous):

in order to solve it im pretty sure u have to log it first

OpenStudy (nincompoop):

no

OpenStudy (anonymous):

so you dont natural log it?

OpenStudy (nincompoop):

it is up to you, this is where all of your algebra knowledge come into play

OpenStudy (nincompoop):

your goal is to achieve a non-zero denominator

OpenStudy (anonymous):

i do remember you can switch lim and ln around and put lim on outside and ln on inside. i did that and i dont know how to follow up with it. :( im trying to achieve the non zero denominator...

OpenStudy (nincompoop):

it doesn't matter how crazy your numerator is. your issue will be your denominator being zero

OpenStudy (nincompoop):

try whatever you're talking about and we'll see how is that going to pan out

OpenStudy (anonymous):

\[\lim_{x \rightarrow 6}(\ln (\sqrt[5]{x^2-4}-2) - \ln (x-6))\]

OpenStudy (anonymous):

is where i got to.... after using my method

OpenStudy (anonymous):

i dont know where to go after that

OpenStudy (nincompoop):

can you rewrite your question because it is not displaying correctly

OpenStudy (anonymous):

This limit problem is a perfect candidate for L'Hopital's Rule.

OpenStudy (anonymous):

That explains why I don't get it... our teacher hasn't covered this part yet. Thanks!

OpenStudy (nincompoop):

how come the problem is not rendering correctly on my screen...?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 6}(\frac{ \sqrt[5]{x^2-4}-2 }{ x-6 })\]

OpenStudy (anonymous):

^thats the problem if u still want to see it

OpenStudy (nincompoop):

show the hospital's way (laughing out loud) @Easyaspi314

OpenStudy (nincompoop):

I am happy to learn this if it didn't give us a DNE

OpenStudy (nincompoop):

limit DNE

OpenStudy (anonymous):

Sure...The derivative of the denominator is just 1...so we just need the derivative of numerator...and that is (using chain rule), (2x/5)(x^2 - 4)^(-4/5)......as x----> 6, that's just 3/20.

OpenStudy (anonymous):



OpenStudy (anonymous):

And it's quite evident from the graph...that it approaches .15 or 3/20.

OpenStudy (nincompoop):

I graphed it and it gave me asymptotes

OpenStudy (anonymous):

Sorry, then you didn't enter it correctly. Be careful as you need parenthesis..plenty of them. I showed a pretty accurate graph of the function.

OpenStudy (anonymous):

Enter the function as: ((x^2 - 4)^(.2) - 2)/(x-6)

OpenStudy (anonymous):

That should you give the correct graph.

OpenStudy (nincompoop):

nvm I did a typo I put -2 instead of -4

OpenStudy (anonymous):

ok.

OpenStudy (nincompoop):

it's why I asked you to do l'hospital earlier because the graph I had would not have any limit

OpenStudy (nincompoop):

it is not rendering properly on my screen

OpenStudy (anonymous):

I dont see that on my screen. I see it nicely "typeset".

OpenStudy (nincompoop):

thanks for your help

OpenStudy (anonymous):

no problem. take care.